# Leaf schemes and Hodge loci [^1] February 25, 2025

### v1

@@ 4.2 Bernoulli numbers
See [BS66, 384-386] or [IR90, Chapter 15] for a proof. The articles [Ser73, SD73, Kat73b] use the above theorem to built up the theory of p-adic modular forms.
@@
This proposition follows from a statement in [SD73, page 19]: “if $f$ is a modular form and $A$ the additive group generated by the coefficients of the $q$-series expansion of $f$, then $f$ has a unique expression as an isobaric element of $A[Q,\Delta]\oplus RA[Q,\Delta]$”. Here, $Q=E_{4},R=E_{6}$ and $\Delta=\frac{1}{1728}(Q^{3}-R^{2})$. The proof is easy, even though at first it did not appear to me and Frederico Bianchini reminded me the argument. First, $f-f_{0}E_{4}^{a}E_{6}^{b}$, for some $a,b\in\mathbb{N}$ such that $k:=4a+6b$ is the weight of $f$, is a cusp form. We know that the ideal of cusp form over $\mathbb{Q}$, is generated by $\Delta$ and so $\frac{f-f_{0}E_{4}^{a}E_{6}^{b}}{\Delta}$ has coefficients in $A$ and is a modular form of weight $k-12$. The proof is finished by using induction on $k$. ∎
@@
${{\cal I}}_{p}:=\left\{P\in\mathbb{F}_{p}[t_{1},t_{2},t_{3}]\ \ \Big{|}\ \ P(a_{1}E_{2},a_{2}E_{4},a_{3}E_{6})=0\right\}.$
This is an ideal in $\mathbb{F}_{p}[t_{1},t_{2},t_{3}]$ and we know that it has many elements. We have $E_{p-1}=1+\frac{2(p-1)}{B_{p-1}}(q+\cdots)$ and Von Staut Clausen theorem says that ${\rm ord}_{p}\frac{2(p-1)}{B_{p-1}}=+1$ and so $E_{p-1}\equiv_{p}1$. We write $E_{p-1}=A(a_{2}E_{4},a_{3}E_{6})$ and by 4.1 the prime $p$ does not appear in the denominator of $A$ and so it makes sense to consider $A\in\mathbb{F}_{p}[t_{2},t_{2}]$. We get $A-1\in{{\cal I}}_{p}$. In a similar way we use Kummer’s theorem for $(p-1)\not|(p+1)$ and so ${\rm ord}_{p}\frac{B_{p+1}}{2(p+1)}\geq 0$ and this is $\equiv_{p}\frac{B_{2}}{4}=24$ which implies that
(49) ${\rm ord}_{p}\frac{B_{p+1}}{2(p+1)}=0$
and hence it it invertible modulo $p$. This together with Fermat’s little theorem imply that $E_{p+1}\equiv E_{2}$. By 4.1 and (49) we can write $E_{p+1}\equiv_{p}B(a_{2}E_{4},a_{6}E_{6})$ with $B\in\mathbb{F}_{p}[t_{2},t_{3}]$, and we have another element $B+12t_{1}\in{{\cal I}}_{p}$.
@@
    It is invariant under the Ramanujan vector field ${\sf v}$ and $({\sf v}^{p}-{\sf v})\mathbb{F}_{p}[t_{1},t_{2},t_{3}]\subset{{\cal I}}_{p}$.
@@
    The scheme ${\rm Zero}({{\cal I}}_{p})$ is an irreducible curve in $\mathbb{A}_{\bar{\mathbb{F}}_{p}}^{3}$ with the smooth point $a$ and tangent to the vector $b$ , both defined in 4.3, at $a$.
@@
1. Since $\frac{1}{12}B(t_{2},t_{3})+t_{1}\in{{\cal I}}_{p}$, we need to show that any $P(t_{2},t_{3})\in{{\cal I}}_{p}$ is a multiple of $A(t_{2},t_{3})-1$. This has been proved in [SD73, Theorem 2 (iv), page 22], see also [Ser73, page 196].
@@ 2. Let us consider the following map which is a ring homomorphism:
In ${\mathfrak{R}}[[q]]$ we consider the derivation $-q\frac{\partial}{\partial q}$ and it turns out that the following is commutative:
(50) $\begin{array}[]{ccc}{\mathfrak{R}}[t_{1},t_{2},t_{3}]&\stackrel{{\scriptstyle{\sf v}}}{{\to}}&{\mathfrak{R}}[t_{1},t_{2},t_{3}]\\ \downarrow&&\downarrow\\ {\mathfrak{R}}[[q]]&\stackrel{{\scriptstyle-q\frac{\partial}{\partial q}}}{{\to}}&{\mathfrak{R}}[[q]].\end{array}$
This implies that ${{\cal I}}_{p}$ is invariant under the Ramanujan vector field ${\sf v}$. One can compute ${\sf v}(A-1)$ and ${\sf v}(B+12t_{1})$ in terms of $A-1$ and $B+12t_{1}$ using the equalities in [SD73, Theorem 2 page 22]. By Fermat little theorem we have $a^{p}\equiv_{p}a$ for all $a\in\mathbb{N}$, and so, $(-q\frac{\partial}{\partial q})^{p}=-q\frac{\partial}{\partial q}$. This implies that ${\sf v}^{p}f-{\sf v}f\in{{\cal I}}_{p}$ for all $f\in{\mathfrak{R}}[t]$.
@@ 3. Note that
$\left\{P\in\bar{\mathbb{F}}_{p}[t_{1},t_{2},t_{3}]\ \ \Big{|}\ \ P(E_{2},E_{4},E_{6})=0\right\}={{\cal I}}_{p}\otimes_{\mathbb{F}_{p}}\bar{\mathbb{F}}_{p}.$
This follows by considering the same equality for the vector space of polynomials $P$ of degree $\leq d$ and the fact that the left hand side of the above equality is defined over $\mathbb{F}_{p}$. By definition ${{\cal I}}_{p}\otimes_{\mathbb{F}_{p}}\bar{\mathbb{F}}_{p}$ is a prime ideal. The point $a$ is a singular point of the Ramanujan vector field. Of course we can prove this also by explicit generators $A-1$ and $B+12t_{1}$ of ${{\cal I}}_{p}$. The smoothness at $a$ follows from $(p-1)A=4t_{2}\frac{\partial A}{\partial t_{2}}+6t_{3}\frac{\partial A}{\partial t_{3}}$. ∎
@@
We remark that ${{\cal I}}_{p}$ is not generated by $({\sf v}^{p}-{\sf v})t_{i},\ \ i=1,2,3$. For the computation below we use the Ramanujan vector field corresponding with the solution $(E_{2},E_{4},E_{6})$ (without constant $a_{i}$’ s). We first compute the linear part of ${\sf v}$ at ${t}$. In the cordinates $x_{i}=t_{i}-1,\ \ i=1,2,3$, ${\sf v}$ can be written as
${\sf v}:=\begin{bmatrix}\frac{\partial}{\partial x_{1}}&\frac{\partial}{\partial x_{2}}&\frac{\partial}{\partial x_{3}}\end{bmatrix}\begin{bmatrix}\frac{1}{6}&\frac{-1}{12}&0\\ \frac{1}{3}&\frac{1}{3}&\frac{-1}{3}\\ \frac{1}{2}&-1&\frac{1}{2}\end{bmatrix}\begin{bmatrix}x_{1}\\ x_{2}\\ x_{3}\end{bmatrix}+\cdots$
@@
$J=\begin{bmatrix}0&1&0\\ 0&0&0\\ 0&0&1\end{bmatrix}\ \ \ S=\begin{bmatrix}1/3&2&1/21\\ 2/3&0&-10/21\\ 1&0&1\end{bmatrix}$
This implies that the linear part of ${\sf v}^{p}-{\sf v}$ is of rank one. If $({\sf v}^{p}-{\sf v})t_{i},\ \ i=1,2,3$ generate ${{\cal I}}_{p}$, this must be $2$ because ${\rm Zero}({{\cal I}}_{p})$ is smooth at $t=(1,1,1)$.
@@ 4.3 Hasse-Witt invariant
For the definition of Hasse-Witt invariants, we follow [AH19]. We actually use a notion of Hasse-Witt invariant for differential forms of the second kind. For a curve $X$ defined over a perfect field of characteristic $p$, there exists a unique map $C:\Omega^{1}_{X}\to\Omega^{1}_{X}$, called the Cartier operator, such that
@@
This operator induces a $1/p$-linear map on merormorphic differential 1-forms on $X$. We can compute the Cartier operator in the following way. Let $a$ be a closed point of a smooth curve $X$ and let $t$ be a coordinate system at $a$. The ${\cal O}_{X,a}^{p}$-module ${\cal O}_{X,a}$ is freely generated by functions $1,t,\cdots,t^{p-1}$. Any meromorphic 1-form which is holomorphic at $a$ admits an expression
(51) $\omega=(\sum_{j=0}^{p-1}f_{j}^{p}t^{j})dt,\ \ f_{j}\in{\cal O}_{X,a}.$
@@
We follow [AH19] and compute the matrix of $C\frac{x^{i}dx}{y}$. Let $f(x)=4x^{3}-t_{2}x-t_{3}$, where $t_{2},t_{3}\in{\mathfrak{k}}$ and ${\mathfrak{k}}$ is a perfect field of characteristic $p$. We denote the inverse of the Frobenius map ${\mathfrak{k}}\to{\mathfrak{k}},\ \ t\mapsto t^{p}$ by $t\mapsto t^{\frac{1}{p}}$ and write
@@
By the $\mathbb{G}_{m}$ action $x\to kx,\ t_{2}\to k^{-2}t_{2},\ t_{3}\to k^{-3}t_{3}$, we can see that $c_{i}\in\mathbb{Z}[t_{2},t_{3}],\ \deg(t_{3})=6,\ \deg(t_{2})=4$ is homogeneous of degree $3(p-1)-2i$. A simple calculation shows that
@@
$[C(\frac{dx}{y}),C(\frac{xdx}{y})]=[\alpha,\omega]\begin{bmatrix}c_{p-1}^{\frac{1}{p}}&c_{p-2}^{\frac{1}{p}}\\ 0&0\end{bmatrix}.$
@@
A hint for the proof of the first congruence can be found in [SD73, last paragraph of page 23]. In this article we read ”This may be proved in one of two ways. On the one hand Deligne has shown that the $q$-series expansion of the Hasse invariant reduces to 1; and Theorem 2 shows that this property characterizes $A$ among polynomials of weight $l-1$. On the other hand the differential equation derived from (ii) is just that which the Hasse invariant is known to satisfy, see Igusa.” The first proof is reproduced in [Kat73b, page 90], see also [Ser73, Theorem 3]. The second congruency at first seemed to be a novelty which has not deserved the attention of masters. The author consulted F. Voloch regarding this which resulted in the following comments. Serre in a note (Algèbre et géométrie, page 81) mentions the article [Rob80] in which a multiplication by $E_{p+1}$ map is characterized uniquely with a certain property related to Hecke operators. This is also reproduced in [Edi92, Proposition 7.2]. At no point its relation with the Cartier map and the differential form of the second kind $\frac{xdx}{y}$ is discussed. From another perspective, Katz in [Kat77, page 57] describes the action of Frobenius map (dual to Cartier) on the de Rham cohomology of elliptic curves. He uses the letter $B$ to denote a coefficient which must be essentially $c_{p-2}$, but no relation with $E_{p+1}$ is discussed. After a personal communication with N. Katz, the author came to know that the puzzle has been only solved many decades later in [Kat21, Theorem 3.1]. This might partially justify the author’s rediscovery of the second congruency. The reader who dislikes proofs using some complicated language of Algebraic Geometry and prefers experimental verification, might use the following computer code in Singular which verifies the proposition for as much as primes that the computer can handle. Here is the code for primes $\leq 300$.
@@
(of 4.2) The curve we are looking for in 4.2 is the one explicitly described in 4.2. Let $(t_{1},t_{2},t_{3})$ be a ${\mathfrak{k}}$-rational point of ${\rm Zero}({{\cal I}}_{p})$, and hence, $A(t_{2},t_{3})=1,\ B(t_{2},t_{3})=t_{1}$. We write the family of elliptic curves in 4.1 in the affine coordinates and write it in the form $y^{2}=4x^{3}-t_{2}x-t_{3},\ \alpha=\frac{dx}{y},\ \omega=\frac{(x+t_{1})dx}{y}$. By 4.3 we know that
$C(\alpha)=A^{\frac{1}{p}}\alpha=\alpha,$
@@
(of 4.3) We have an algebraic curve $C$ in ${\sf T}/\mathbb{F}_{p}$ and its smooth point $a$ in the discriminant loci $\Delta=0$ and we know that it is tangent to ${\sf v}$. Moreover, we have the vector $b$ given in 4.3 tangent to $C$ at $a$. Since $C$ is smooth at $a$, we can parameterize it, that is, there are formal power series $t_{i}:=\sum_{i=0}^{\infty}t_{i,n}q^{n}\in\mathbb{F}_{p}[[q]],\ \ i=1,2,3$ such that $(t_{1,0},t_{2,0},t_{3,0})=a$ and $(t_{1,1},t_{2,1},t_{3,1})=b$. For $t=(t_{1},t_{2},t_{3})$, both formal power series $\frac{\partial t}{\partial q}$ and ${v}(t(q))$ are tangent to the curve $C$ and the first one evaluated at $a$ is non-zero. Therefore, we have ${\sf v}(t)=a(q)\frac{\partial t}{\partial q}$, where $a(q)=\sum_{i=0}^{\infty}a_{i}q^{i}\in\mathbb{F}_{p}[[q]]$. Since $\frac{\partial t}{\partial q}(a)=b\not=0$ and ${\sf v}(a)=0$, we conclude that $a_{0}=0$. Moreover, we compute the coefficient of $q$ in ${\sf v}(t(q))$ and see it is $-b$, that is,
@@
see [Mov21a, Page 333]. This implies that $a_{1}=-1$. By inverse function theorem (or formal change of coordinates tangent to the identity) we can assume that $a(q)=-q$. Therefore, $q\frac{\partial t}{\partial q}={\sf v}(t)$ and this implies the $p$-integrality of the solution over $\mathbb{Q}$. ∎
@@
(of 4.4) We claim that the $B+12t_{1}A$ is the first integral of ${\sf v}$, that is, ${\sf v}(B+12t_{1}A)=0$. For this we use the differential equations of $A$ and $B$ in [SD73, Theorem 2 (ii), page 22], see also [Ser73]. Under the transformation $(t_{1},t_{2},t_{3})\to(kt_{1},k^{2}t_{2},k^{3}t_{3})$ with $k\in{\mathfrak{k}}$, ${\sf v}$ is mapped to $k^{-2}{\sf v}$ and the curve $C_{1}:A=1,\ B=t_{1}$ is mapped to another curve $C_{k}$ passing through $(ka_{1},k^{2}a_{2},k^{3}a_{3})$. All these curves lie in the hypersurface $B+12t_{1}A=0$ which intersects the discriminant hypersurface $\Delta=0$ at two components ${\rm Sing}({\sf v})$ and $t_{2}=t_{3}=0$. This shows that ${\sf v}$ restricted to $B+12t_{1}A$ has the first integral $A=-\frac{B}{12t_{1}}$. ∎
@@
(52) ${\sf v}:=y\frac{\partial}{\partial x}+\frac{1}{2}P^{\prime}(x)\frac{\partial}{\partial y}$
is a derivation/vector field on $E$. Over complex numbers, the Weierstrass uniformization $z\mapsto[\wp(z):\wp^{\prime}(z):1]$ maps $\frac{\partial}{\partial z}$ to ${\sf v}$. For the vector field (52), it is known that

### v2

@@ 4.2 Bernoulli numbers
See [BorevichShafarevich, 384-386] or [IrelandRosen, Chapter 15] for a proof. The articles [Serre1973, Swinnerton1973, ka73] use the above theorem to build up the theory of p-adic modular forms. Recall the Eisenstein series:
(4.3) $E_{2k}(\tau)=1-\frac{4k}{B_{2k}}\sum_{n=1}^{\infty}\sigma_{2k-1}(n)q^{n},\ \ k=1,2,3,\ldots.$
where $\sigma_{i}(n):=\sum_{d\mid n}d^{i}$.
@@
This proposition follows from a statement in [Swinnerton1973, page 19]: “if $f$ is a modular form and $A$ the additive group generated by the coefficients of the $q$-series expansion of $f$, then $f$ has a unique expression as an isobaric element of $A[Q,\Delta]\oplus RA[Q,\Delta]$”. Here, $Q=E_{4},R=E_{6}$ and $\Delta=\frac{1}{1728}(Q^{3}-R^{2})$. The proof is easy, even though at first it did not appear to me and Frederico Bianchini reminded me the argument. First, $f-f_{0}E_{4}^{a}E_{6}^{b}$, for some $a,b\in\mathbb{N}$ such that $k:=4a+6b$ is the weight of $f$, is a cusp form. We know that the ideal of cusp form over $\mathbb{Q}$, is generated by $\Delta$ and so $\frac{f-f_{0}E_{4}^{a}E_{6}^{b}}{\Delta}$ has coefficients in $A$ and is a modular form of weight k12k-12. The proof is finished by using induction on $k$. ∎
@@
${{\mathcal{I}}}_{p}:=\left\{P\in\mathbb{F}_{p}[t_{1},t_{2},t_{3}]\ \ \Big|\ \ P(a_{1}E_{2},a_{2}E_{4},a_{3}E_{6})=0\right\}.$
This is an ideal in $\mathbb{F}_{p}[t_{1},t_{2},t_{3}]$ and we know that it has many elements. We have $E_{p-1}=1+\frac{2(p-1)}{B_{p-1}}(q+\cdots)$ and Von Staut Clausen theorem says that ${\rm ord}_{p}\frac{2(p-1)}{B_{p-1}}=+1$ and so $E_{p-1}\equiv_{p}1$. We write
(4.4) $E_{p-1}=A\left(\frac{1}{12}E_{4},-\frac{1}{216}E_{6}\right)$
[…]
(4.5) ${\rm ord}_{p}\frac{B_{p+1}}{2(p+1)}=0$
and hence it it invertible modulo $p$. This together with Fermat’s little theorem imply that $E_{p+1}\equiv E_{2}$. By 4.1 and (4.5) we can write
(4.6) $E_{p+1}\equiv_{p}B\left(\frac{1}{12}E_{4},-\frac{1}{216}E_{6}\right)$
with $B\in\mathbb{F}_{p}[t_{2},t_{3}]$, and we have another element $B+12t_{1}\in{{\mathcal{I}}}_{p}$.
@@
    It is invariant under the Ramanujan vector field $\sf v$} and $({\sf v}^{p}-{\sf v})\mathbb{F}_{p}[t_{1},t_{2},t_{3}]\subset{{\mathcal{I}}}_{p}$.
@@
    The scheme ${\rm Zero}({{\mathcal{I}}}_{p})$ is an irreducible curve in $\mathbb{A}_{\bar{\mathbb{F}}_{p}}^{3}$ with the smooth point $a$ and tangent to the vector $b$ at $a$, both defined in 4.3.
@@
1. Since $\frac{1}{12}B(t_{2},t_{3})+t_{1}\in{{\mathcal{I}}}_{p}$, we need to show that any $P(t_{2},t_{3})\in{{\mathcal{I}}}_{p}$ is a multiple of $A(t_{2},t_{3})-1$. This has been proved in [Swinnerton1973, Theorem 2 (iv), page 22], see also [Serre1973, page 196].
@@ 2. Let us consider the following map which is a ring homomorphism:
In ${\mathfrak{R}}[[q]]$ we consider the derivation $q\frac{\partial}{\partial q}$ and it turns out that the following is commutative:
(4.7) $\begin{array}[]{ccc}{\mathfrak{R}}[t_{1},t_{2},t_{3}]&\stackrel{{\scriptstyle{\sf v}}}{{\to}}&{\mathfrak{R}}[t_{1},t_{2},t_{3}]\\ \downarrow&&\downarrow\\ {\mathfrak{R}}[[q]]&\stackrel{{\scriptstyle-q\frac{\partial}{\partial q}}}{{\to}}&{\mathfrak{R}}[[q]].\end{array}$
This implies that ${{\mathcal{I}}}_{p}$ is invariant under the Ramanujan vector field $\sf v$}. One can compute ${\sf v}(A-1)$ and ${\sf v}(B+12t_{1})$ in terms of A1A-1 and $B+12t_{1}$ using the equalities in [Swinnerton1973, Theorem 2 page 22]. By Fermat little theorem we have $a^{p}\equiv_{p}a$ for all $a\in\mathbb{N}$, and so, $(-q\frac{\partial}{\partial q})^{p}=-q\frac{\partial}{\partial q}$. This implies that ${\sf v}^{p}f-{\sf v}f\in{{\mathcal{I}}}_{p}$ for all $f\in{\mathfrak{R}}[t]$.
@@ 3. Note that
$\left\{P\in\bar{\mathbb{F}}_{p}[t_{1},t_{2},t_{3}]\ \ \Big|\ \ P(E_{2},E_{4},E_{6})=0\right\}={{\mathcal{I}}}_{p}\otimes_{\mathbb{F}_{p}}\bar{\mathbb{F}}_{p}.$
This follows by considering the same equality for the vector space of polynomials $P$ of degree $\leq d$ and the fact that the left hand side of the above equality is defined over $\mathbb{F}_{p}$. By definition ${{\mathcal{I}}}_{p}\otimes_{\mathbb{F}_{p}}\bar{\mathbb{F}}_{p}$ is a prime ideal. The point $a$ is a singular point of the Ramanujan vector field. Of course we can prove this also by explicit generators A1A-1 and $B+12t_{1}$ of ${{\mathcal{I}}}_{p}$. The smoothness at $a$ follows from $(p-1)A=4t_{2}\frac{\partial A}{\partial t_{2}}+6t_{3}\frac{\partial A}{\partial t_{3}}$. ∎
@@
We remark that ${{\mathcal{I}}}_{p}$ is not generated by $({\sf v}^{p}-{\sf v})t_{i},\ \ i=1,2,3$. For the computation below we use the Ramanujan vector field corresponding with the solution $(E_{2},E_{4},E_{6})$ (without constant $a_{i}$’ s). We first compute the linear part of $\sf v$} at t{t}. In the cordinates $x_{i}=t_{i}-1,\ \ i=1,2,3$, $\sf v$} can be written as
${\sf v}:=\begin{bmatrix}\frac{\partial}{\partial x_{1}}&\frac{\partial}{\partial x_{2}}&\frac{\partial}{\partial x_{3}}\end{bmatrix}\begin{bmatrix}\frac{1}{6}&\frac{-1}{12}&0\\ \frac{1}{3}&\frac{1}{3}&\frac{-1}{3}\\ \frac{1}{2}&-1&\frac{1}{2}\end{bmatrix}\begin{bmatrix}x_{1}\\ x_{2}\\ x_{3}\end{bmatrix}$$\cdots$
@@
$J=\begin{bmatrix}0&1&0\\ 0&0&0\\ 0&0&1\end{bmatrix}\ \ \ S=\begin{bmatrix}1/3&2&1/21\\ 2/3&0&-10/21\\ 1&0&1\end{bmatrix}$.
This implies that the linear part of ${\sf v}^{p}-{\sf v}$ is of rank one. If $({\sf v}^{p}-{\sf v})t_{i},\ \ i=1,2,3$ generates ${{\mathcal{I}}}_{p}$, this must be $2$ because ${\rm Zero}({{\mathcal{I}}}_{p})$ is smooth at $t=(1,1,1)$.
@@ 4.3 Hasse-Witt invariant
For the definition of Hasse-Witt invariants, we follow [AchterHowe]. We actually use a notion of Hasse-Witt invariant for differential forms of the second kind. For a curve $X$ defined over a perfect field of characteristic $p$, there exists a unique map $C:\Omega^{1}_{X}\to\Omega^{1}_{X}$, called the Cartier operator, such that
@@
This operator induces a $1/p$-linear map on merormorphic differential 1-forms on $X$. We can compute the Cartier operator in the following way. Let $a$ be a closed point of a smooth curve $X$ and let $t$ be a coordinate system at $a$. The ${\mathcal{O}}_{X,a}^{p}$-module ${\mathcal{O}}_{X,a}$ is freely generated by functions $1,t,\cdots,t^{p-1}$. Any meromorphic 1-form which is holomorphic at $a$ admits an expression
(4.8) $\omega=(\sum_{j=0}^{p-1}f_{j}^{p}t^{j})dt,\ \ f_{j}\in{\mathcal{O}}_{X,a}.$
@@
We follow [AchterHowe] and compute the matrix of $C\frac{x^{i}dx}{y}$. Let $f(x)=4x^{3}-t_{2}x-t_{3}$, where $t_{2},t_{3}\in{\mathfrak{k}}$ and $\mathfrak{k}$} is a perfect field of characteristic $p$. We denote the inverse of the Frobenius map ${\mathfrak{k}}\to{\mathfrak{k}},\ \ t\mapsto t^{p}$ by $t\mapsto t^{\frac{1}{p}}$ and write
@@
By the $\mathbb{G}_{m}$ action $x\to kx,\ t_{2}\to k^{-2}t_{2},\ t_{3}\to k^{-3}t_{3}$, we can see that $c_{i}\in\mathbb{Z}[t_{2},t_{3}],\ \deg(t_{3})=6,\ \deg(t_{2})=4$ is homogeneous of degree 3(p1)2i3(p-1)-2i. A simple calculation shows that
@@
$\left[C(\frac{dx}{y}),C(\frac{xdx}{y})\right]=\left[\frac{dx}{y},\frac{xdx}{y}\right]\begin{bmatrix}c_{p-1}^{\frac{1}{p}}&c_{p-2}^{\frac{1}{p}}\\ 0&0\end{bmatrix}$.
@@
A hint for the proof of the first congruence can be found in [Swinnerton1973, last paragraph of page 23]. In this article we read ”This may be proved in one of two ways. On the one hand Deligne has shown that the $q$-series expansion of the Hasse invariant reduces to 1; and Theorem 2 shows that this property characterizes $A$ among polynomials of weight l1l-1. On the other hand the differential equation derived from (ii) is just that which the Hasse invariant is known to satisfy, see Igusa.” The first proof is reproduced in [ka73, page 90], see also [Serre1973, Theorem 3]. The second congruency at first seemed to be a novelty which has not deserved the attention of masters. The author consulted F. Voloch regarding this which resulted in the following comments. Serre in a note (Algèbre et géométrie, page 81) mentions the article [Robert1980] in which a multiplication by $E_{p+1}$ map is characterized uniquely with a certain property related to Hecke operators. This is also reproduced in [Edixhoven1992, Proposition 7.2]. At no point its relation with the Cartier map and the differential form of the second kind $\frac{xdx}{y}$ is discussed. From another perspective, Katz in [Katz1977, page 57] describes the action of Frobenius map (dual to Cartier) on the de Rham cohomology of elliptic curves. He uses the letter $B$ to denote a coefficient which must be essentially $c_{p-2}$, but no relation with $E_{p+1}$ is discussed. After a personal communication with N. Katz, the author came to know that the puzzle has been only solved many decades later in [Katz2021, Theorem 3.1]. This might partially justify the author’s rediscovery of the second congruency. The reader who dislikes proofs using some complicated language of Algebraic Geometry and prefers experimental verification, might use the following computer code in Singular which verifies the proposition for as much as primes that the computer can handle. Here is the code for primes $\leq 300$.
@@
(of 4.2) The curve we are looking for in 4.2 is the one explicitly described in 4.2. Let $(t_{1},t_{2},t_{3})$ be a $\mathfrak{k}$}-rational point of ${\rm Zero}({{\mathcal{I}}}_{p})$, and hence, $A(t_{2},t_{3})=1,\ -\frac{1}{12}B(t_{2},t_{3})=t_{1}$. We write the family of elliptic curves in 4.1 in the affine coordinates and write it in the form $y^{2}=4x^{3}-t_{2}x-t_{3},\ \alpha=\frac{dx}{y},\ \omega=\frac{(x+t_{1})dx}{y}$. By 4.3 we know that
$C(\alpha)=A^{\frac{1}{p}}\alpha=\alpha,\ \$
@@
(of 4.3) 1) implies 2): We have an algebraic curve $C={\rm Zero}(I)$ in ${\sf T}/\mathbb{F}_{p}$ and its smooth point $a$ in the discriminant loci $\Delta=0$ and we know that it is tangent to $\sf v$}, that is, ${\sf v}(I)\subset I$. Moreover, we have the vector $b$ given in 4.3 tangent to $C$ at $a$. Since $C$ is smooth at $a$, we can parameterize it, that is, there are formal power series $\check{t}_{i}(q):=\sum_{i=0}^{\infty}\check{t}_{i,n}q^{n}\in\mathbb{F}_{p}[[q]],\ \ i=1,2,3$ such that
$(\check{t}_{1,0},\check{t}_{2,0},\check{t}_{3,0})=a,\ (\check{t}_{1,1},\check{t}_{2,1},\check{t}_{3,1})=b$
and any element in the ideal $I$ of $C$ annihilates $\check{t}(q):=(\check{t}_{1}(q),\check{t}_{2}(q),\check{t}_{3}(q))$. Both vectors $\frac{\partial\check{t}(q)}{\partial q}$ and ${\sf v}(\check{t}(q)):=({\sf v}_{1}(\check{t}(q)),\ {\sf v}_{2}(\check{t}(q)),{\sf v}_{3}(\check{t}(q)))$, where ${\sf v}=\sum_{i=1}^{3}{\sf v}_{i}\frac{\partial}{\partial t_{i}}$ are tangent to the curve $C$, that is,
$\sum_{i=1}^{3}A_{i}\frac{\partial f}{\partial t_{i}}=0,\ \forall f\in I,$
for $A$ one of these vectors. Since the constant term of $\frac{\partial\check{t}}{\partial q}$ is the non-zero vector $b$, we have
${\sf v}(\check{t}(q))=a(q)\frac{\partial\check{t}}{\partial q},\hbox{ where }a(q)=\sum_{i=0}^{\infty}a_{i}q^{i}\in\mathbb{F}_{p}[[q]].$
Moreover, from $\frac{\partial\check{t}}{\partial q}(a)=b\not=0$ and ${\sf v}(a)=0$, we conclude that $a_{0}=0$. We compute the coefficient of $q$ in ${\sf v}(\check{t}(q))$ and see that it is b-b, that is,
@@
see [ho13, Page 333]. This implies that $a_{1}=-1$. We claim that there is a formal change of variables $q\mapsto\tilde{q}:=f(q)\in\mathbb{F}_{p}[[q]]$ which maps the derivation $a(q)\frac{\partial}{\partial q}$ to $\tilde{q}\frac{\partial}{\partial\tilde{q}}$. For this we use ${\sf v}^{p}={\sf v}$ restricted to the curve $C$ and we restate and prove it in 4.4. If this is the case, we can assume that a(q)=qa(q)=-q. Therefore, $q\frac{\partial t}{\partial q}={\sf v}(t)$ and this implies the $p$-integrality of the solution over $\mathbb{Q}$.
1) implies 2): We use the fact that $t_{i}=a_{i}E_{2i}$ and the curve $C$ is given by $A-1=B+12t_{1}A=0$. ∎
Let $\mathbb{A}^{1,{\rm for}}_{\mathbb{F}_{p}}$ be the spectrum of $\mathbb{F}_{p}[[q]]$.
Proposition 4.4.
 [^1] This proposition is a particular case of the so called Poincaré linearization theorem in characteristic $p$, see 6.9.
Let ${\sf v}=a(q)\frac{\partial}{\partial q}=(\lambda q+\cdots)\frac{\partial}{\partial q},\lambda\in\mathbb{F}_{p}$ be a vector field in $\mathbb{A}^{1,{\rm for}}_{\mathbb{F}_{p}}$. There is a coordinate system $\tilde{q}$ in $\mathbb{A}^{1,{\rm for}}_{\mathbb{F}_{p}}$ such that ${\sf v}=\lambda\tilde{q}\frac{\partial}{\partial\tilde{q}}$ if and only if ${\sf v}^{p}={\sf v}$.
@@
The direction $\Rightarrow$ is easy: Let $\tilde{q}=f(q):=q+\cdots\in\mathbb{F}_{p}[[q]]$. The map
$\mathbb{A}^{1,{\rm for}}_{\mathbb{F}_{p}}\to\mathbb{A}^{1,{\rm for}}_{\mathbb{F}_{p}},\ q\mapsto\tilde{q}$,
maps $a(q)\frac{\partial}{\partial q}$ to $\lambda\tilde{q}\frac{\partial}{\partial\tilde{q}}$ and the the second vector field satisfies ${\sf v}^{p}={\sf v}$ and hence the first vector field satisfies too. Note that we use Fermat’s little theorem $\lambda^{p}=\lambda$.
For $\Leftarrow$ we argue as follows: We have to find $f(q):=q+\cdots\in\mathbb{F}_{p}[[q]]$ such that ${\sf v}(f)=\lambda f$. We claim that
$f=-\sum_{i=1}^{p-1}(\lambda^{-1}{\sf v})^{i}q$
satisfies this. The equality ${\sf v}(f)=\lambda f$ follows from $(\lambda^{-1}{\sf v})^{p}=(\lambda^{-1}{\sf v})$. Moreover, the coefficient of $q^{1}$ in $P$ is $1=-(1+1+\cdots+1)$, (p1)(p-1)-times. ∎
Proof.
(of 4.4) We claim that the $B+12t_{1}A$ is the first integral of $\sf v$}, that is, ${\sf v}(B+12t_{1}A)=0$. For this we use the differential equations of $A$ and $B$ in [Swinnerton1973, Theorem 2 (ii), page 22], see also [Serre1973]. Under the transformation $(t_{1},t_{2},t_{3})\to(kt_{1},k^{2}t_{2},k^{3}t_{3})$ with $k\in{\mathfrak{k}}$, $\sf v$} is mapped to $k^{-2}{\sf v}$ and the curve $C_{1}:A=1,\ B=t_{1}$ is mapped to another curve $C_{k}$ passing through $(ka_{1},k^{2}a_{2},k^{3}a_{3})$. All these curves lie in the hypersurface $B+12t_{1}A=0$ which intersects the discriminant hypersurface $\Delta=0$ at two components ${\rm Sing}({\sf v})$ and $t_{2}=t_{3}=0$. This shows that $\sf v$} restricted to $B+12t_{1}A$ has the first integral $A=-\frac{B}{12t_{1}}$. ∎
@@
(4.9) ${\sf v}:=y\frac{\partial}{\partial x}+\frac{1}{2}P^{\prime}(x)\frac{\partial}{\partial y}$
is a derivation/vector field on $E$. Over complex numbers, the Weierstrass uniformization $z\mapsto[\wp(z):\wp^{\prime}(z):1]$ maps $\frac{\partial}{\partial z}$ to $\sf v$}. For the vector field (4.9), it is known that